OpenStudy (anonymous):

the trapezoids are similar. the are of the smaller one is 564 m^2. find the area of the larger trapezoid to the nearest whole number. medal. thanks :)

OpenStudy (amistre64):

without knowing a scalar factor to compare the similar shapes by, this would prove most difficult

OpenStudy (amistre64):

the area of a trap is the average base time height

OpenStudy (anonymous):

lol, wow.. forgot to draw it. give me a second

OpenStudy (amistre64):

pfft, i added the height instead of multiplying

OpenStudy (amistre64):

\[\frac{sb_1+sb_2}{2}sh\] \[s^2(\frac{b_1+b_2}{2}h)\]

OpenStudy (amistre64):

24*s = 57, find s

OpenStudy (anonymous):

never mind, got it. :) but thanks :) @amistre64